# JWST Observations of Starbursts: PAHs Closely Trace the Cool Phase of M82’s Galactic Wind

### v1

@@ 3.3 PAH Intensity, Gas Mass, and the Radiation Field
The analyses in Sections 3.1 and 3.2 show that PAH emission correlates strongly with the CO(1–0) and H emission near the disk and moderately so in the outflow. This suggests that PAHs are physically associated with the cool gas of the wind. However, the nearly fixed PAH-to-CO ratio seen in Figure 3 and 4 is unexpected. For stochastically heated grains, the intensity of PAH emission is expected to depend on local physical conditions as
@@
where is the intensity of the interstellar radiation field, the column density of dust-bearing gaseous material, the dust-to-gas ratio, and is the abundance of PAH molecules relative to the overall dust mass (Draine & Li, 2007, and see discussion in Leroy et al. 2023, Sandstrom et al. 2023, Chown et al. 2025). We lack constraints for and in the outflow, but and change with elevation (Leroy et al., 2015).
@@
Figure 5 explores this behavior. The top row shows the profiles of PAH intensity (here Spitzer m444We use Spitzer as it closely matches F770W and covers the full region, minimizing edge effects.) along with surface densities of H2 traced by CO(1–0), atomic gas (HI), and the total neutral gas surface density (), as a function of distance along the minor axis. The second row shows the PAH emission normalized by the molecular, atomic, and total gas column density as a function of distance from the starburst. The final row shows three estimates of the intensity of the radiation field: from PDR modeling (Levy et al., 2023), from IR SED modeling (Leroy et al., 2015) (with models from Draine & Li 2007), and one assuming , where is the radial distance from the galaxy center. Following Eq. 2, we expect .
@@
The mild decline in the PAH-gas ratio is consistent with the decline of the profile over the same range. Thus the observations appear consistent with Equation 2, with the PAHs mixed with all neutral gas components and responding to heating by the interstellar radiation field. We do not have independent constraints on and 555Our adopted from Leroy et al. (2015) assumes a fixed over the range of our study.. The simulations of Richie & Schneider (2025) find that decreases with distance from the starburst due to dust destruction. Any decrease in (or , e.g., due to selective PAH destruction) would decrease at fixed . Given the good match between and in the bottom panel of Figure 5, such variations do not seem to be required by the data out to kpc.
4 Conclusions
@@ Appendix A The Effect of Vertical Distance Scaling on PAH-Gas Phase Correlations
We find that when correcting for vertical distance (as discussed in Section 3.1), only the CO(1–0) and H datasets still show a moderate correlation with the F770W data past 2 kpc. The X-ray data does not present any correlation (Spearman rank coefficient of near zero), showing that the dust and hot gas are decoupled at these large distances and only correlate near the starburst. We find near identical results for the F1130W (not shown).
@@ Appendix B H Interface Zoom In
As hypothesized by Lopez et al. (2025), the H emission may trace turbulent mixing layers modeled in multiphase galactic wind simulations (Tan & Fielding, 2024), which can cool rapidly and prolong cloud survival. As shown in the main text, the strong correlation between molecular gas and PAH emission demonstrates that JWST images provide a high-resolution proxy for cold gas. Thus, Figure B1 likely captures the interfaces between PAH-traced cold clouds and the hot wind, where H emission arises from their surfaces, presumably due to shocks.

### v2

@@ III.3 PAH Intensity, Gas Mass, and the Radiation Field
The analyses in Sections III.1 and III.2 show that PAH emission correlates strongly with the CO(1–0) and H emission near the disk and moderately so in the outflow. This suggests that PAHs are physically associated with the cool gas of the wind. However, the nearly fixed PAH-to-CO ratio seen in Figure 3 and 4 is unexpected. For stochastically heated grains, the intensity of PAH emission is expected to depend on local physical conditions as
@@
where is the intensity of the interstellar radiation field, the column density of dust-bearing gaseous material, the dust-to-gas ratio, and is the abundance of PAH molecules relative to the overall dust mass (Draine and Li, 2007, and see discussion in Leroy et al. 2023, Sandstrom et al. 2023, Chown et al. 2025). We lack constraints for and in the outflow, but and change with elevation (Leroy et al., 2015).
@@
Figure 5 explores this behavior. The top row shows the profiles of PAH intensity (here Spitzer m444We use Spitzer as it closely matches F770W and covers the full region, minimizing edge effects.) along with surface densities of H2 traced by CO(1–0), atomic gas (HI), and the total neutral gas surface density (), as a function of distance along the minor axis. We have convolved all the data in this section to 27′′ and reproject onto a shared astrometric grid. The second row shows the PAH emission normalized by the molecular, atomic, and total gas column density as a function of distance from the starburst. The final row shows three estimates of the intensity of the radiation field: from PDR modeling (Levy et al., 2023), from IR SED modeling (Leroy et al., 2015) (with models from Draine and Li 2007), and one assuming , where is the radial distance from the galaxy center. Following Eq. 2, we expect .
@@
The mild decline in the PAH-gas ratio is consistent with the decline of the profile over the same range. Thus the observations appear consistent with Equation 2, with the PAHs mixed with all neutral gas components and responding to heating by the interstellar radiation field. We do not have independent constraints on and 555Our adopted from Leroy et al. (2015) assumes a fixed over the range of our study.. The simulations of Richie and Schneider (2025) find that decreases with distance from the starburst due to dust destruction. Any decrease in (or , e.g., due to selective PAH destruction) would decrease at fixed . Given the good match between and in the bottom panel of Figure 5, such variations do not seem to be required by the data out to kpc.
IV Conclusions
@@ Appendix A The Effect of Vertical Distance Scaling on PAH-Gas Phase Correlations
We find that when correcting for vertical distance (as discussed in Section III.1), only the CO(1–0) and H datasets still show a moderate correlation with the F770W data past 2 kpc. The X-ray data does not present any correlation (Spearman rank coefficient of near zero), showing that the dust and hot gas are decoupled at these large distances and only correlate near the starburst. We find near identical results for the F1130W (not shown).
@@ Appendix B H Interface Zoom In
As hypothesized by Lopez et al. (2025), the H emission may trace turbulent mixing layers modeled in multiphase galactic wind simulations (Tan and Fielding, 2024), which can cool rapidly and prolong cloud survival. As shown in the main text, the strong correlation between molecular gas and PAH emission demonstrates that JWST images provide a high-resolution proxy for cold gas. Thus, Figure B1 likely captures the interfaces between PAH-traced cold clouds and the hot wind, where H emission arises from their surfaces, presumably due to shocks.